Suppose a two sound source emitted a slightly different sound frequency, 262Hz ans 266 Hz, were released at the same time. Determine the beat frequency and frequency heard by the observer.


What are the given;

First Frequency (F1)    F1= 262 Hz

Second Frequency (F2)  F2= 266 Hz


What are the missinv:  

Beat Frequency       Fb= ?

Perceived Frequency   Fo=?


1
Expert's answer
2021-02-17T11:01:02-0500

Beat Frequency


"f_B=266-262=4\ Hz"

Perceived Frequency


"f_o=0.5(266+262)=264\ Hz"